I know this isn't on this certian subject but it was the rang of questions i could find. I'm having trouble learning/understanding indirect variations. Can someone expain it for me. here is a problem: y=9, x=1, x=?, and y=10. Please help me with this problem!! thanks
kayla

Answer by stanbon(48535)

If x and y are indirectly related they satisfy the formula y=k/x
You are told that y=9 if x=1.
That allows us to determine what "k" is, as follows:
9=k/1 or k=9
We now know the relevant formula is y=9/x
So you have the question: "What is "x" if y=10?"
Put those values into the equation and solve for "x", as follows:
10=9/x
10x=9
So x=9/10 when y=10.
Hope this helps.
Cheers,
stan H.